OpenStudy (anonymous):

how do you solve log base 2 of 1/2 = x ?

OpenStudy (anonymous):

\[\log_2(\frac{1}{2})=x\iff 2^x=\frac{1}{2}\]

OpenStudy (anonymous):

and since \[2^{-1}=\frac{1}{2}\], -1 is your answer

OpenStudy (anonymous):

okay thank you !